# A Pre-trained Sequential Recommendation Framework: Popularity Dynamics for Zero-shot Transfer

### 03 Jan 2024 v1

@@ 4. PrepRec Framework
Here, we first introduce the model architecture of PrepRec (§ 4.1) and the training procedure (§ 4.2). Then, we formally define the zero-shot inference process (§ 4.3).
@@ 4.1. Model Architecture
First, we transform the user sequence $\{v_{u,1},v_{u,2},...,v_{u,|\mathcal{S}_{u}|}\}$ for each user $u$ into a fixed-length sequence $\mathcal{S}_{u}$ = $\{v_{u,1},v_{u,2},...,v_{u,L}\}$ via truncating or padding, where $L$ is a pre-defined hyper-parameter controlling the maximum length of the sequence. Note that the truncating is done by removing the oldest items in the sequence.
4.1.1. Universal Item Representation encoder
Previous works in sequential and pre-trained sequential recommender systems learn to represent items either through learnable embeddings specific to the item IDs, or through the item’s textual information. However, item IDs are not generalizable across domains and textual information is not always available and is application-dependent. In this work, we propose a novel item representation encoder that learns to represent items through the changes in their popularity histories over different time periods (popularity dynamics).
Formally, given an $v_{j}$ that has interaction in time period $t$, denoted as $v_{j}^{t}$, we define two popularity representations for $v_{j}^{t}$: popularity $\mathbf{p}_{j}^{t}\in\mathbb{R}^{k}$ over a coarse time period (e.g., month) and popularity $\mathbf{h}_{j}^{t}\in\mathbb{R}^{k}$ over a fine time period (e.g., week). To calculate $\mathbf{p}_{j}^{t}$ and $\mathbf{h}_{j}^{t}$ , we first count the number of interactions of $v_{j}^{t}$ over the two time period, denoted as $a_{j}^{t}\in\mathbb{R}$ (coarse period number of interactions) and $b_{j}^{t}\in\mathbb{R}$ (fine period number of interactions). Specifically, we calculate them as:
@@ 4.1.1. Universal Item Representation encoder
where $\gamma\in\mathbb{R}$ is the discount factor and $c_{a}(v_{j}^{m})$ is the number of interactions of $v_{j}$ over a coarse time period $m$. Similarly, $c_{b}(v_{j}^{t})$ denotes the number of interactions of $v_{j}$ over a fine time period $t$. Note that we do not have a discounting factor when computing $b_{j}^{t}$ since we want it to capture the current popularity information, whereas $a_{j}^{t}$ captures the cumulative popularity of an item. We then calculate $a_{j}^{t}$ and $b_{j}^{t}$’s percentiles relative to their corresponding coarser and finer popularity distributions over all items at time $t$ and denote this by $P(a_{j}^{t})\in\mathbb{R}$ and $P({b_{j}^{t}})\in\mathbb{R}$, respectively. Finally, we encode $P(a_{j}^{t})$ and $P({b_{j}^{t}})$ into $k$ dimension vector representations $\mathbf{p}_{j}^{t}$ and $\mathbf{h}_{j}^{t}$ respectively. The encoding details can be found in Appendix A.
We now define the popularity dynamics of $v_{j}$ at time $t$ over the coarse time period (long-term horizon) to be $\{\mathbf{p}_{j}^{1},\mathbf{p}_{j}^{2},...,\mathbf{p}_{j}^{t-1}\}$, and over the fine time period (short-term horizon) as $\{\mathbf{h}_{j}^{1},\mathbf{h}_{j}^{2},...,\mathbf{h}_{j}^{t-1}\}$. The reason we use $t-1$ is to constrain access to future interactions and prevent information leakage, i.e., we do not have access to the popularity statistics of $v_{j}$ at time $t$ if we are at time $t$. For example, say an interaction happens on the second Wednesday in February, we consider the coarser and finer time period up until the end of January and the end of the first week in February respectively. To limit computation, we consider window sizes $m,n$ for $\mathcal{P}$ and $\mathcal{H}$ respectively. Formally, we now define the coarse popularity dynamics of $v_{j}$ at time $t$ as $\mathcal{P}_{j}^{t}=\{\mathbf{p}_{j}^{t-m},\mathbf{p}_{j}^{t-m+1},...,\mathbf{p}_{j}^{t-1}\}$, and the fine popularity dynamics of $v_{j}$ at time $t$ as $\mathcal{H}_{j}^{t}=\{\mathbf{h}_{j}^{t-n},\mathbf{h}_{j}^{t-n+1},...,\mathbf{h}_{j}^{t-1}\}$.
We then compute the embedding of item $v_{j}$ at time $t$ via the universal item representation encoder, defined as a function $\mathcal{E}(\mathcal{P}_{j}^{t},\mathcal{H}_{j}^{t})$ that learns to encode the popularity dynamics $\mathcal{P}_{j}^{t}$ and $\mathcal{H}_{j}^{t}$ into a $d$ dimension vector representation $\mathbf{e}_{j}^{t}$. Specifically, we have:
@@
4.1.2. Relative Time Interval Encoding
Since we explicitly consider time when modeling sequences, we propose a relative time interval encoding to encode the time intervals between consecutive interactions of a user. Formally, given an interaction sequence $\mathcal{S}_{u}=\{v_{u,1},v_{u,2},...,v_{u,L}\}$ of user $u$, we define the time interval between $v_{u,j}$ and $v_{u,j+1}$ as $t_{u,j}=t(v_{u,j+1})-t(v_{u,j})$, where $t(v_{u,j})$ is the time that user $u$ interacts with item $v_{u,j}$. We then rank the time intervals of user $u$. Define the rank of relative time interval of $t_{u,j}$ as $r_{u,j}=\text{rank}(t_{u,j})$. The relative time interval encoding of interval $t_{u,j}$ is then defined as ${\bm{T}}_{r_{u,j}}\in\mathbb{R}^{d}$, where ${\bm{T}}\in\mathbb{R}^{L\times d}$, following the same setup in (Vaswani et al., 2017), is a fixed sinusoidal encoding matrix defined as:
@@ 4.1.2. Relative Time Interval Encoding
We also tried a learnable time interval encoding, but this didn’t perform as well as the sinusoidal encoding. We hypothesize that the sinusoidal encoding is more generalizable across domains and the learnable encoding is more prone to overfitting.
@@ 4.1.3. Positional Encoding
4.1.4. Popularity Dynamics-Aware Transformer Layer
We follow previous works in sequential recommendation (Kang and McAuley, 2018; Sun et al., 2019; Li et al., 2020) and propose an extension to the self-attention mechanism for modeling the popularity dynamics of items in a sequence by incorporating model components introduced in § 4.1.1§ 4.1.2, and § 4.1.3.
Given a user sequence $\mathcal{S}_{u}=\{v_{u,1},v_{u,2},...,v_{u,L}\}$, we first compute in input matrix ${\bm{E}}_{u}$ as:
@@ 4.1.4. Popularity Dynamics-Aware Transformer Layer
Causality: In sequential recommendation, the prediction of the $s+1$ item that a user will interact with should only depend on the first $s$ items that the user has interacted with in the past. However, the $s$-th output of the multi-head self-attention layer contains all the input information. Therefore, as in (Li et al., 2020; Kang and McAuley, 2018), we do not let the model attend to the future items by forbidding links between $Q_{i}$ and $K_{j}$ ($j>i$) in the attention function.
Point-Wise Feed-Forward Network: While the multi-head self-attention mechanism can capture the dependencies between items in a sequence and incorporate all previous items’ information, it is still inherently a linear model. Therefore, to add nonlinearity and interactions between different embedding dimensions, we follow previous works in sequential recommendation (Sun et al., 2019; Kang and McAuley, 2018; Li et al., 2020) and apply the same point-wise feed-forward network to the output of each multi-head self-attention layer. Formally, suppose the output of the multi-head self-attention layer is $\textbf{z}_{u}$, the point-wise feed-forward network is defined as:
@@ 4.1.5. Prediction
Given a sequence $\mathcal{S}_{u}$ of user $u$, we denote the embedding of user $u$ after going through § 4.1.1 § 4.1.2 § 4.1.3 § 4.1.4 as $\mathbf{q}_{u}$ (Figure 1). Suppose at time $t^{+}$, we want to compute user $u$’s preference of item $v_{j}$, we first compute the item representation $\mathbf{e}_{j}^{t^{+}}$ from § 4.1.1. Then, we predict the score as the innner product of $\mathbf{q}_{u}$ and $\mathbf{e}_{j}^{t^{+}}$, formally:
@@
There is no information leakage in the prediction process, i.e., we do not assume access to the popularity statistics of $v_{j}$ at time $t^{+}$ if we are at time $t^{+}$.
@@ 4.2. Training Procedure
Now we show how to train the PrepRec model. Similar to (Kang and McAuley, 2018), we adopt the binary cross entropy loss as the objective function, formally:
@@
where $\mathcal{S}_{u,:z}\coloneqq\{v_{u,1},v_{u,2},...,v_{u,z}\}$. $v^{t}_{z+1}$ represents the $z+1$-th item in the sequence that happened at time $t$. We use Adam (Kingma and Ba, 2014) as the optimizer and train the model end-to-end. Note that compared to previous sequential recommenders, we do not have any parameters specific to the item IDs. Essentially, we are only optimizing ${\bm{W}}_{p}$ and parameters related to the multi-head self-attention mechanism.
@@ 4.3. Zero-shot Inference
Our inference setup is defined in § 3. For zero-shot transfer, we first compute two levels of popularity dynamics on the target dataset. Then, we directly apply the model pre-trained on the source dataset to get the prediction.
Formally, suppose we are given a pre-trained model $\mathcal{F}$ trained on ${\bm{M}}$. $\mathcal{F}$ is essentially the scoring function learned from source domain ${\bm{M}}$. Let us denote the interaction matrix of the target domain as ${\bm{M}}^{\prime}$. We first compute the popularity dynamics of each item in ${\bm{M}}^{\prime}$ over a coarser time period and a finer time period. Then, we apply the pre-trained model $\mathcal{F}$ to ${\bm{M}}^{\prime}$ and compute the prediction score as:
@@
Note that in this procedure, we use the pre-trained model $\mathcal{F}$ trained on domain ${\bm{M}}^{\prime}$ to predict the next item $v^{t}_{j^{\prime}}$ that user $u^{\prime}$ will interact with in domain ${\bm{M}}^{\prime}$. We do not use any auxiliary information in both domains. In addition, none of the parameters in $\mathcal{F}$ are updated during the inference process.
To summarize, in this section, we showed how to develop a pre-trained sequential recommender system based on the popularity dynamics of items. We enforce the structure of each interaction in the sequence by the positional encoding and introduce a relative time encoding for modeling time intervals between two consecutive interactions. In addition, we showed the training process and formally defined the zero-shot inference procedure. In the next section, we will evaluate PrepRec on multiple datasets.
@@ 5. Experiments
Table 3. Regular sequential recommendation results. We embolden the best results and mark the best baseline results with $*{}^{\prime}*^{\prime}$. Interp represents the interpolation results between PrepRec and BERT4Rec. PrepRec $\Delta$ denotes the performance difference between PrepRec and the best results among the selected baselines, similar for Interp $\Delta$. We want to emphasize that the regular sequential recommendation is not the main focus of this paper, rather, we include the results here to show the importance of modeling popularity dynamics in the sequential recommendation and establish a baseline of PrepRec’s performance. Despite not focusing on the performance of regular sequential recommendation, PrepRec achieves comparable performance to the state-of-the-art sequential recommenders, with only on average 0.2% worse than the best performing sequential recommenders in R@10 while having only a fraction of the model size (Table 9). After a simple post-hoc interpolation, we outperform the state-of-the-art sequential recommenders by 13.8% in R@10 on average.
We present extensive experiments on five real-world datasets to evaluate the performance of PrepRec, following the problem settings in § 3. We introduce the following research questions (RQ) to guide our experiments: (RQ1) How well can PrepRec perform under the regular, in-domain sequential recommendation setting? This helps us understand the importance of modeling popularity dynamics in sequential recommendation. (RQ2) Can PrepRec (popularity dynamics) generalize to zero-shot cross-domain and cross-application transfer? (RQ3) What are the factors that affect the performance of PrepRec ? In other words, how sensitive is PrepRec to different choices in constructing popularity dynamics? (RQ4) What is the effect of the amount of training data available for PrepRec ? This can be crucial if the pre-trained model is applied to a new domain with limited data.
5.1. Datasets and Preprocessing
We evaluate our proposed method on five real-world datasets, with varying applications, sizes, and density levels.
The dataset statistics are presented in Table 2. We compute the density as the ratio of the number of interactions to the number of users times the number of items. The Tool dataset is the largest in terms of the number of users and items, with Office and Movie following, and Epinions and Music smallest. Both datasets from Douban (i.e., movie and music) are the densest, while the Amazon review datasets (i.e., office and tool) are the sparsest.
For fair evaluation, we follow the same preprocessing procedure as previous works (Kang and McAuley, 2018; Sun et al., 2019; Rendle et al., 2010; Tang and Wang, 2018), i.e., we binarize the explicit ratings to implicit feedback. In addition, for each user, we sort interactions by their timestamp and take these chronological sequences as input, using the second most recent action for validation, and the most recent action for testing. As described in § 4.1, our method directly utilizes interaction timestamps to discover long-term and short-term popularity trends. Note that existing sequential methods  (Sun et al., 2019; Kang and McAuley, 2018) use timestamps for sequence construction, so our method isn’t any more restrictive in application.

### 10 Apr 2024 v2

@@ 4. PrepRec Framework
We first introduce the model architecture of PrepRec (§ 4.1) then the training procedure (§ 4.2). Finally, we formally define the zero-shot inference process (§ 4.3).
@@ 4.1. Model Architecture
The first step of building a pre-trained sequential recommender is to learn universal item representations. Prior works in sequential and pre-trained sequential recommender systems learn to represent items either through learnable embeddings specific to their item IDs or through their textual metadata, largely limiting the generalizability of the model under various scenarios, e.g., different domains, platforms, and applications. Our solution is to exploit the item popularity statistics available in every domain to learn universal item representations. Instead of explicit item IDs or auxiliary information, we learn to represent items at a given timestamp through the changes in their popularity histories over different periods (popularity dynamics). We propose a popularity dynamics-aware Transformer architecture that obtains the representation of users’ behavior sequences through item popularity dynamics.
4.1.1. Item Popularity Encoder
We learn to represent items based on their popularity dynamics, i.e., changes in their popularity histories. Intuitively, we should model items’ popularity by two horizons: long-term and short-term. Long-term horizons reflect the overall popularity of items, whereas short-term horizons should capture the recent trends in the domain. For example, the long-term popularity of a winter coat measures how popular is the coat in general, while its short-term popularity reflects more temporal changes, e.g., season, weather conditions, and fashion trends. Therefore, consider an item $v_{j}$ that has interaction at time $t$, denoted as $v_{j}^{t}$, we define two popularity representations for $v_{j}^{t}$: popularity $\mathbf{p}_{j}^{t}\in\mathbb{R}^{k}$ over a coarse period (e.g., month) and popularity $\mathbf{h}_{j}^{t}\in\mathbb{R}^{k}$ over a fine period (e.g., week).
To calculate $\mathbf{p}_{j}^{t}$ and $\mathbf{h}_{j}^{t}$, we first calculate the popularities of $v_{j}^{t}$ over the two horizons, denoted as $a_{j}^{t}\in\mathbb{R}^{+}$ (coarse period number of interactions) and $b_{j}^{t}\in\mathbb{R}^{+}$ (fine period number of interactions). Specifically, we calculate them as:
@@ 4.1.1. Item Popularity Encoder
where $\gamma\in\mathbb{R}^{+}$ is a pre-defined discount factor and $c_{a}(v_{j}^{m})$ is the number of interactions of $v_{j}$ over a coarse time period $m$. Similarly, $c_{b}(v_{j}^{t})$ denotes the number of interactions of $v_{j}$ over a fine period $t$. We do not impose the discounting factor when computing $b_{j}^{t}$ since we want it to capture the current popularity information, whereas $a_{j}^{t}$ captures the cumulative popularity of an item over a longer horizon.
The popularity of an item is a real value, which is not comparable across items from different domains. Therefore, we calculate the percentiles of $a_{j}^{t}$ and $b_{j}^{t}$ relative to their corresponding coarser and finer popularity distributions over all items at time $t$, denoted as $P(a_{j}^{t})\in\mathbb{R}^{+}$ and $P({b_{j}^{t}})\in\mathbb{R}^{+}$, respectively.
We now encode the popularity percentiles $P(a_{j}^{t})$ and $P({b_{j}^{t}})$ into $k$ dimensional vector representations $\mathbf{p}_{j}^{t}$ and $\mathbf{h}_{j}^{t}$ respectively. Denote the popularity encoder as $E_{p}:\mathbb{R}^{+}\rightarrow\mathbb{R}^{k}$, which takes in a percentile value. Suppose given the popularity percentile $P(a_{j}^{t})\in\mathbb{R}^{+}$ over a coarse time period $t$, the coarse level popularity vector representation $\mathbf{p}_{j}^{t}\in\mathbb{R}^{k}$ is computed as follows:
$\mathbf{p}_{j}^{t}$ $=E_{p}(P(a_{j}^{t}))$
$(\mathbf{p}_{j}^{t})_{i}$ =$=\begin{cases}1-\{\frac{P}{k-1}\},&\text{if }i=\lfloor\frac{P}{k-1}\rfloor\\ \{\frac{P}{k-1}\},&\text{if }i=\lfloor\frac{P}{k-1}\rfloor+1\\ 0,&\text{otherwise}\end{cases}$
[…]
4.1.2. Universal Item Representation
We now define the popularity dynamics of $v_{j}$ at time $t$ over the coarse period (long-term horizon) to be $\mathcal{P}_{j}^{t}=\{\mathbf{p}_{j}^{1},\mathbf{p}_{j}^{2},...,\mathbf{p}_{j}^{t-1}\}$, and over the fine period (short-term horizon) as $\mathcal{H}_{j}^{t}=\{\mathbf{h}_{j}^{1},\mathbf{h}_{j}^{2},...,\mathbf{h}_{j}^{t-1}\}$. We use $t-1$ to constrain access to future interactions and prevent information leakage, i.e., we do not have access to the popularity statistics of $v_{j}$ at time $t$ if we are at time $t$. For example, say an interaction happens on the second Wednesday in February, we consider the coarser and finer time period up until the end of January and the end of the first week in February respectively. To limit computation, we constrain window sizes $m,n$ for $\mathcal{P}$ and $\mathcal{H}$ respectively. Formally, the coarse popularity dynamics of $v_{j}$ at time $t$ is $\mathcal{P}_{j}^{t}=\{\mathbf{p}_{j}^{t-m},\mathbf{p}_{j}^{t-m+1},...,\mathbf{p}_{j}^{t-1}\}$, and the fine popularity dynamics of $v_{j}$ at time $t$ is $\mathcal{H}_{j}^{t}=\{\mathbf{h}_{j}^{t-n},\mathbf{h}_{j}^{t-n+1},...,\mathbf{h}_{j}^{t-1}\}$.
Finally, we compute the embedding of item $v_{j}$ at time $t$ via the universal item representation encoder, defined as a function $\mathcal{E}(\mathcal{P}_{j}^{t},\mathcal{H}_{j}^{t})$ that learns to encode the popularity dynamics $\mathcal{P}_{j}^{t}$ and $\mathcal{H}_{j}^{t}$ into a $d$ dimension vector representation $\mathbf{e}_{j}^{t}$. Specifically, we have:
@@ 4.1.2. Universal Item Representation
4.1.3. Relative Time Interval
We also consider the time interval between two consecutive interactions when modeling sequences. Differences in time intervals might indicate differences in the users’ behaviors. While previous works explore absolute time intervals (Li et al., 2020), different domains exhibit diverse time scales, thus making modeling absolute time intervals ungeneralizable. Therefore, we propose to encode relative time intervals into modeling sequences. Given an interaction sequence $\mathcal{S}_{u}=\{v_{u,1},v_{u,2},...,v_{u,L}\}$ of user $u$, we define the time interval between $v_{u,j}$ and $v_{u,j+1}$ as $t_{u,j}=t(v_{u,j+1})-t(v_{u,j})$, where $t(v_{u,j})$ is the time that user $u$ interacts with item $v_{u,j}$. We then rank the time intervals of user $u$. Define the rank of relative time interval of $t_{u,j}$ as $r_{u,j}=\text{rank}(t_{u,j})$. The relative time interval encoding of interval $t_{u,j}$ is then defined as ${\bm{T}}_{r_{u,j}}\in\mathbb{R}^{d}$, where ${\bm{T}}\in\mathbb{R}^{L\times d}$, following the same setup in (Vaswani et al., 2017), is a fixed sinusoidal encoding matrix defined as:
@@ 4.1.3. Relative Time Interval
We also tried a learnable time interval encoding, but it yielded worse performance. We hypothesize that the sinusoidal encoding is more generalizable across domains and the learnable encoding is more prone to overfitting.
@@ 4.1.4. Positional Encoding
4.1.5. Popularity Dynamics-Aware Transformer
We follow previous works in sequential recommendation (Kang and McAuley, 2018; Sun et al., 2019; Li et al., 2020) and propose an extension to the self-attention mechanism by incorporating universal item representations (§ 4.1.2), relative time intervals(§ 4.1.3), and positional encoding (§ 4.1.4).
Firstly, we transform the user sequence $\{v_{u,1},v_{u,2},...,v_{u,|\mathcal{S}_{u}|}\}$ for each user $u$ into a fixed-length sequence $\mathcal{S}_{u}$ = $\{v_{u,1},v_{u,2},...,v_{u,L}\}$ via truncating the oldest interactions or padding, where $L$ is a pre-defined hyper-parameter controlling the maximum length of the sequence. Given a user sequence $\mathcal{S}_{u}=\{v_{u,1},v_{u,2},...,v_{u,L}\}$, we compute its input matrix ${\bm{E}}_{u}$ as:
@@ 4.1.5. Popularity Dynamics-Aware Transformer
Causality: In sequential recommendation, the prediction of the $L+1$ item should only depend on the first $L$ items that the user has interacted with in the past. However, the $L$-th output of the multi-head self-attention layer contains all the input information. Therefore, as in (Li et al., 2020; Kang and McAuley, 2018), we do not let the model attend to the future items by forbidding links between $Q_{i}$ and $K_{j}$ ($j>i$) in the attention function.
Point-Wise Feed-Forward Network: To add nonlinearity and interactions between different embedding dimensions, we follow previous works in sequential recommendation (Sun et al., 2019; Kang and McAuley, 2018; Li et al., 2020) and apply the same point-wise feed-forward network to the output of each multi-head self-attention layer. Formally, suppose the output of the multi-head self-attention layer is $\textbf{z}_{u}$, the point-wise feed-forward network is defined as:
@@ 4.1.6. Prediction
Given a sequence $\mathcal{S}_{u}$ of user $u$ as input, we denote $\mathbf{q}_{u}$ as the output of the popularity dynamics-aware transformer. Suppose at time $t^{+}$, we want to predict the likelihood of $v_{j}$ being the next item in the sequence, we first compute the item representation $\mathbf{e}_{j}^{t^{+}}$ from § 4.1.2. Then, we predict the score as the inner product of $\mathbf{q}_{u}$ and $\mathbf{e}_{j}^{t^{+}}$, formally:
@@
Note that there is no information leakage in the prediction process, i.e., we do not assume access to the popularity statistics of $v_{j}$ at time $t^{+}$ if we are at time $t^{+}$ (§ 4.1.2).
@@ 4.2. Training Procedure
Now we present how to train the PrepRec model. Similar to (Kang and McAuley, 2018), we adopt the binary cross entropy loss as the objective function, formally:
@@
where $\mathcal{S}_{u,:z}\coloneqq\{v_{u,1},v_{u,2},...,v_{u,z}\}$. $v^{t}_{z+1}$ represents the $z+1$-th item in the sequence that happened at time $t$. We use Adam (Kingma and Ba, 2014) as the optimizer and train the model end-to-end. Note that compared to previous sequential recommenders, we do not have any parameters modeling item IDs. Essentially, we are only optimizing ${\bm{W}}_{p}$ and parameters related to the multi-head self-attention mechanism.
@@ 4.3. Zero-shot Inference
Suppose we are given a pre-trained model $\mathcal{F}$ trained on ${\bm{M}}$, where $\mathcal{F}$ is the scoring function learned from source domain ${\bm{M}}$. Denote the interaction matrix of the target domain as ${\bm{M}}^{\prime}$. We first compute the popularity dynamics of each item in ${\bm{M}}^{\prime}$ over a coarser period and a finer period. Then, we apply the pre-trained model $\mathcal{F}$ to ${\bm{M}}^{\prime}$ and compute the prediction score as:
@@
Note that in this procedure, we use the pre-trained model $\mathcal{F}$ trained on domain ${\bm{M}}^{\prime}$ to predict the next item $v^{t}_{j^{\prime}}$ that user $u^{\prime}$ will interact with in domain ${\bm{M}}^{\prime}$. We do not use any auxiliary information in either domain. In addition, none of the parameters in $\mathcal{F}$ are updated during the zero-shot inference process.
To summarize, in this section, we showed how to develop a pre-trained sequential recommender system based on the popularity dynamics of items. We enforce the structure of each interaction in the sequence by the positional encoding and introduce a relative time encoding for modeling time intervals between two consecutive interactions. In addition, we showed the training process and formally defined the zero-shot inference procedure. In the next section, we present experiments to evaluate PrepRec .
@@ 5. Experiments
We present extensive experiments on five real-world datasets to evaluate the performance of PrepRec, following the problem settings in § 3. We introduce the following research questions (RQ) to guide our experiments: (RQ1) How well can PrepRec perform under the regular, in-domain sequential recommendation setting? (RQ2) Can PrepRec (popularity dynamics) generalize to zero-shot cross-domain and cross-application transfer? (RQ3) How different model components affect the performance of PrepRec ? (RQ4) What is the difference between the predictions of PrepRec and those of the state-of-the-art sequential recommenders?
Dataset Office Tool Movie Music Epinions
SasRec
1.331m
3.581m
2.044m
0.542m
1.054m
BERT4Rec
2.687m
7.233m
4.126m
1.094m
2.127m
TiSasRec
1.367m
3.617m
2.127m
0.578m
1.09m
PrepRec
0.045m
0.045m
0.045m
0.045m
0.045m
Table 2. Comparison of model sizes (i.e., number of learnable parameters in millions) over different datasets. PrepRec uses significantly fewer number of parameters than baselines, ranging from 12 to 90x smaller.
5.1. Datasets and Preprocessing
We evaluate our proposed method on five real-world datasets across different applications, with varying sizes, and density levels.
We present dataset statistics in Table 1. We compute the density as the ratio of the number of interactions to the number of users times the number of items. Both datasets from Douban (i.e., movie and music) are the densest and have no auxiliary information available, while the Amazon review datasets (i.e., office and tool) are the sparsest.
For fair evaluation, we follow the same preprocessing procedure as previous works (Kang and McAuley, 2018; Sun et al., 2019; Rendle et al., 2010; Tang and Wang, 2018), i.e., we binarize the explicit ratings to implicit feedback. In addition, for each user, we sort interactions by their timestamp and take these chronological sequences as input, using the second most recent action for validation, and the most recent action for testing.
@@ 5.1. Datasets and Preprocessing
Table 3. Regular sequential recommendation results. We embolden the best results and mark the best baseline results with $*{}^{\prime}*^{\prime}$. Interp represents the interpolation results between PrepRec and BERT4Rec. PrepRec $\Delta$ denotes the performance difference between PrepRec and the best results among the selected baselines, similar for Interp $\Delta$. Despite not focusing on the performance of regular sequential recommendation, PrepRec achieves comparable performance to the state-of-the-art sequential recommenders, with only on average 0.2% worse than the best performing sequential recommenders in R@10 while having only a fraction of the model size (Table 2). After a simple post-hoc interpolation, we outperform the state-of-the-art sequential recommenders by 13.8% in R@10 on average.
5.2. Baselines
To assess the effectiveness of our proposed method, we compare against the following set of baselines. These baselines include classic general recommendation models and state-of-the-art sequential recommendation models.
  • BPR (Rendle et al., 2009) learns latent representations for users and items through matrix factorization.
  • MArank (Yu et al., 2019) captures individual and union-level item dependency through multi-order attention.
To the best of our knowledge, we are the first to tackle the zero-shot sequential recommendation without any auxiliary information. (Hou et al., 2022, 2023; Wang et al., 2023b) are designed for few-shot adaptation and uses auxiliary information. (Ding et al., 2021) is capable of zero-shot recommendation but requires textual descriptions of items, which we do not consider in the experiments. Therefore, we do not include the existing pre-trained sequential recommenders as baselines.
5.3. Experimental Setup
Following previous works (Koren, 2008; He et al., 2017; Kang and McAuley, 2018; Sun et al., 2019), we adopt the leave-one-out evaluation method: for each user, we pair its test item with 100 unobserved items according to the user’s interaction history. Then we rank the test item for the user among the 101 total items. We use two standard evaluation metrics for top-$k$ recommendation: Recall@k (R@k) and Normalized Discounted Cumulative Gain@N (N@k). R@k measures the proportion of test items that appear in the top-$k$ recommendations, and N@k accounts for the ranking position of the test interaction. Our model explicitly utilize popularity information. Therefore, in addition to randomly sampled unobserved items (negatives), we also present results where we sample the negatives based on their popularities, i.e., popular items have higher probabilities of being sampled as negatives, to show that PrepRec can learn discriminative item representations. We report the average of R@k and N@k over all the test interactions.
We use publicly available Pytorch implementations for the baselines. For fair evaluation, we set dimension size $d$ to 50, max sequence length $L$ to $200$, and batch size to $128$ in all models. We use three layers (one base and two additional) in NCF and NGCF, and two attention layers for the sequential models (SasRec, BERT4Rec, TiSasRec, and ours). We use an Adam optimizer and tune the learning rate in the range $\{10^{-4},10^{-3},10^{-2}\}$ and set the weight decay to $10^{-5}$. We use the dropout regularization rate of $0.3$ for all models. We set $\gamma=0.5$ in Equation 1, whose reason we will discuss the reason in  § 5.8.1. We define the coarse and fine period to be 10 and 2 days respectively, and we fix the window size to be $m=12$ and $n=4$ for all datasets (§ 4.1.2). We train PrepRec for a maximum of 80 epochs. All experiments are conducted on a Tesla V100 using PyTorch. We repeat each experiment 5 times with different random seeds and report the average performance.
Dataset Music Office Epinions
Metric
R@10
N@10
R@10
N@10
R@10
N@10
MostPop
0.197
0.139
0.099
0.046
0.163
0.110
SasRec (Kang and McAuley, 2018)
0.749
0.519
0.453
0.291
0.658
0.442
BERT4Rec (Sun et al., 2019)
0.747
0.519
0.461 0.299
0.655
0.456
PrepRec
0.739
0.523
0.443
0.280
0.762 0.551
PrepRec $\Delta$
-1.3%
+0.7%
-2.2%
-6.3%
+15.8%
+17.2%
Table 4. Regular sequential recommendation results with popularity-based negative sampling. We present additional in-domain recommendation results with unobserved (negative) items sampled based on their popularity (Sun et al., 2019) to show that PrepRec can learn discriminative universal item and sequence representations despite depending only on popularity statistics.